LABORATORY ABSORPTION SPECTRA OF CARBON MOLECULES FOR COMPARISON WITH DIB DATA

\underline{JOHN P.MAIER}, Department of Chemistry, University of Basel, Klingelbergstrasse 80, CH-4056 Basel, Switzerland.

The gas-phase electronic absorption spectra of several new carbon molecular systems could be observed in the laboratory allowing a direct comparison with astronomical DIB data. These include the non-linear carbon chains, C_{n}H_{4}^{+} n=4,6,8, C_{5}H_{5} isomers which absorb near the 4429 DIB, and large cyclic systems such as C_{18}, all with absorbtions in the visible.